OpenStudy (anonymous):

Simplify: square root of 32 x to the power of 5 y to the power of 4 z end root 4 x to the power of 2 y to the power of 2 square root of 2 x z end root 16 x to the power of 2 y to the power of 2 square root of 2 x z end root 16 x to the power of 2 y to the power of 2 z square root of 2 x end root 4 x to the power of 2 y to the power of 2 z square root of 2 x end root

OpenStudy (camerondoherty):

Is this your equation? http://prntscr.com/46nbvj

OpenStudy (anonymous):

yes

OpenStudy (driftracer305):

k can u tell me what is the power of 2 to get 32?......... like 2 ^? is 32....?

OpenStudy (camerondoherty):

Is this one of your options? http://prntscr.com/46ncu5

OpenStudy (driftracer305):

it says A..... @camerondoherty

OpenStudy (driftracer305):

(-_-?)

OpenStudy (camerondoherty):

huh?

OpenStudy (anonymous):

yes and 5

OpenStudy (driftracer305):

k @Keridavis .... out of 2^5..... how many of the pair of 2's would come out....from the sqrt sign?

OpenStudy (driftracer305):

k lemme explain....

OpenStudy (driftracer305):

\[\sqrt{2} \times \sqrt{2}\times \sqrt{2} \times \sqrt{2} \times \sqrt{2} = 2 \times 2 \sqrt{2}\]

OpenStudy (driftracer305):

\[4\sqrt{2}\]................ u following @Keridavis ?

OpenStudy (driftracer305):

similarly x^5 .... = \[x^2\sqrt{x}\]

OpenStudy (driftracer305):

and \[\sqrt{y^4} = y^2 \]

OpenStudy (driftracer305):

and what u have left in the sqrt sign is \[\sqrt{2xz}\]

OpenStudy (driftracer305):

so to put it all together........ u have 4 and x^2 and y^2 on the outside and on the inside of sqrt sign u have 2 x z.......... so what would that lookk like @Keridavis ????

OpenStudy (anonymous):

A?

OpenStudy (driftracer305):

yes ur correct @Keridavis